Write (12,22,1) as a linear combination of the vectors (0,2,1),(3,0,1),(5,6,0) if possible


1
Expert's answer
2021-10-29T03:00:49-0400

It is possible to write (12,22,1) as a linear combination of the vectors (0,2,1), (3,0,1), (5,6,0):



"\begin{bmatrix}
 12 \\
 22 \\
1
\end{bmatrix}=2\begin{bmatrix}
 0 \\
 2 \\
1
\end{bmatrix}+(-1)\begin{bmatrix}
 3 \\
 0 \\
1
\end{bmatrix}+3\begin{bmatrix}
 5 \\
 6 \\
0
\end{bmatrix}."